OpenStudy (anonymous):

What is the area of a circle with diameter 10 inches?

OpenStudy (anonymous):

that would be pi*(d/2)^2

hero (hero):

Hint: \(A = \pi r^2 \) and \(r = \frac{d}{2}\)

OpenStudy (anonymous):

78.5 inches.

hero (hero):

You're not supposed to just give the answer.

OpenStudy (anonymous):

its ok it was wrong anyway

hero (hero):

lol

hero (hero):

Were you supposed to enter an exact answer or approximate answer? If the answer is supposed to be appropriate, to how many decimal places?

OpenStudy (anonymous):

well i actually wrote this answer a lot diffrent then it was wrote